### v1

# Thompson Sampling-Based Learning and Control for Unknown Dynamic Systems

Abstract
Thompson sampling (TS) is an effective method to explore parametric uncertainties and can therefore be used for active learning-based controller design. However, TS relies on finite parametric representations, which limits its applicability to more general spaces, which are more commonly encountered in control system design. To address this issue, this work pro poses a parameterization method for control law learning using reproducing kernel Hilbert spaces and designs a data-driven active learning control approach. Specifically, the proposed method treats the control law as an element in a function space, allowing the design of control laws without imposing restrictions on the system structure or the form of the controller. A TS framework is proposed in this work to explore potential optimal control laws, and the convergence guarantees are further provided for the learning process. Theoretical analysis shows that the proposed method learns the relationship between control laws and closed-loop performance metrics at an exponential rate, and the upper bound of control regret is also derived. Numerical experiments on controlling unknown nonlinear systems validate the effectiveness of the proposed method.
Index Terms:
Learning-based Control, Thompson Sampling, Active Learning Control
I Introduction
In many practical scenarios, the technical challenges or high costs of accurately modeling system dynamics have driven the development of data-driven control, whose performance heavily depends on the quality of the data [1, 2, 3]. To address this issue, active learning control (ALC) improves control performance by actively designing the data acquisition process or enhances data utilization efficiency by actively exploring the most informative sampling points [4, 5, 6]. Thus, ALC has become a powerful tool for enhancing control performance in complex and uncertain systems, particularly in applications such as robotics [7], industrial automation [8], and autonomous vehicles [9].
Among other data-driven control methods, ALC designs data acquisition and utilization processes to improve learning and control effectiveness. Depending on the learning objectives, ALC can be categorized into three types: model-oriented, policy-oriented (control law-oriented), and dual control. Model-oriented methods aim to obtain more accurate or reliable models [10, 11, 12, 13, 14]. In [15], the controller is chosen by maximizing the Fisher information to actively learn the Koopman operator dynamics of robots. Policy-oriented methods focus on exploring the impact of different strategies (control laws) [16, 17]. A reward-free reinforcement learning method was proposed in [17], aiming to explore the relationship between different strategies and reward functions. In the above two types of methods, control performance during the exploration process is often a secondary consideration or even ignored. Dual control, on the other hand, utilizes existing information to not only experiment with the system to learn about its behavior and improve control in the future but also effectively control the system [18, 19, 20]. In dual control, exploration is usually conducted under the premise that certain closed-loop properties of the system are guaranteed, such as stability [18], robustness [19], and regret [20], etc.
As an active sampling method capable of balancing exploration and exploitation, the TS method has gained significant attention in recent years [21, 22, 23]. In the TS method, the weight of sampling an element is proportional to its posterior probability, which can be updated using various distributions, such as Gaussian distributions [24], Bernoulli distributions [25], and Poisson distributions [26].
Following the principles of ALC and the effectiveness of TS, a number of works have been proposed to design active learning control strategies using TS. Leveraging the success of numerous applications, the classical TS method has been explored for stochastic control problems with parametric uncertainty. Early works considered optimization problems where the unknown parameters are selected from a finite discrete set. For example, the work in [27] addressed scenarios where unknown parameters were chosen from a finite discrete set, and the control law is designed based on the learned parameters. More related studies include applications in learning parameterized Markov decision processes [28], regularized linear optimization problems [29], channel selection [30], and multi-agent games and consensus problems [23]. Later, in [31], the gap between discrete and continuous spaces was bridged, where the unknown parameters were assumed to belong to a continuous real space. Expanding on these advancements, the ability to effectively learn unknown parameters in continuous spaces has made the TS method more widely applicable in the field of control. For instance, the authors of [32] considered the control problem of stochastic differential systems with diffusion processes and unknown parameters. They not only analyzed the regret of the learning process but also discussed the stability of the controller, thereby providing a comprehensive framework for such systems. Furthermore, a method for efficiently designing linear-quadratic regulators for linear systems was discussed in [33], which highlights the versatility of TS in addressing classical control problems. Beyond the limitations of finite discrete parameter spaces [27] and continuous real spaces [31], TS methods have the potential for further development in optimizing controller design within functional spaces. This motivates the design of a user-friendly and easily generalizable ALC method in this work.
In this work, we consider complex and unknown nonlinear systems with almost no prior information, and design an active learning control strategy based on TS that is transferable and generalizable. Specifically, the controller is treated as an element in a function space. Online data is utilized to update the posterior distribution function of the optimal control law within the function space. The controller is periodically updated online according to the TS strategy, which is performed in a general function space and does not rely on prior knowledge of system structures. Therefore, the proposed method’s generalizability and transferability make it easily applicable to different systems. To achieve this goal, however, several challenges need to be addressed. First, unlike optimizing parameters in real spaces, it is more challenging to characterize function spaces and design algorithms for updating the associated probability distribution functions. Second, the control process and learning process are coupled in TS-based ALC methods, which adds extra difficulty to the analysis of learning convergence and control performance. Finally, compared to the potential optimal control law, the learned control method is influenced by the construction of the function space, the posterior distribution update strategy, and the TS sampling algorithm, making the analysis of control regret more complex.
The contributions of this work are summarized as follows:
  1. 1.
    A control law learning method is proposed through parameterizing a Hilbert space. By constructing Hilbert spaces and a convex hull structure, the proposed method can effectively represent complex nonlinear control laws, which further provides theoretical support for control law optimization. Furthermore, the upper bound of performance degradation (compared to the potential optimal control law) caused by the proposed parameterization method is analyzed (Theorem 1).
  2. 2.
    Within the parameterized Hilbert space, a TS-based control law learning method is designed to balance exploration and exploitation. A Bayesian inference method is proposed to update the posterior distribution of the unknown cost function using online sampled data. Based on the obtained posterior distribution, a control law is further selected according to the TS mechanism. As a result, we prove that the cost function learned by the proposed method converges exponentially to a neighborhood of the true cost function (Theorem 2).
  3. 3.
    The control performance of the proposed method is also analyzed. The upper bound of control regret between the proposed method and the potential optimal control law is studied. Specifically, the control regret bound is discussed in terms of the impact of function space parameterization and TS mechanism. Our result (Theorem 3) shows that the control regret consists of a constant part caused by function space parameterization and an exponentially convergent part caused by posterior distribution updates.
The remainder of this work is organized as follows: Section II formulates the considered ALC problem; Section III introduces some useful definitions and lemmas; Section IV presents the main results of this work, including the parameterization of the function space, the learning method based on TS, the learning convergence analysis, and the discussion of the control regret; Section LABEL:sec:NumericalExperiments provides numerical experiments to validate the effectiveness of the proposed methods; and Section V concludes this work and discusses potential future works.
Notation: Random variables are denoted by lowercase letters, and their specific realizations are represented by uppercase letters with subscripts. Sample spaces are indicated by letters like $\mathcal{X,Y}$, and the corresponding $\sigma$-field is denoted using $\Sigma$. For example, let ${\bm{x}}=[x_{1},x_{2},\ldots,x_{n}]^{\rm T}$ be the state of a nonlinear system with samples $X_{1},~{}X_{2},\ldots$, which take values in the space $\mathcal{X}$ with $\sigma$-field $\Sigma_{\bm{x}}$. For a random variable ${\bm{x}}$, the expectation, variance, and covariance are denoted as $\mathbb{E}[{\bm{x}}]$, $\mathbb{V}[{\bm{x}}]$, and $\mathbb{C}[{\bm{x}},{\bm{y}}]$, respectively. The inner product of two vectors ${\bm{x}}$ and ${\bm{y}}$ is denoted as $\langle{\bm{x}},{\bm{y}}\rangle$. The Cartesian product of sets $\mathcal{G}_{1},\mathcal{G}_{2}$ is represented as $\mathcal{G}_{1}\times\mathcal{G}_{2}$. Moreover, we define $\underline{\bm{x}}\otimes\overline{\bm{x}}:=\{{\bm{x}}=[{x}_{1},\ldots,{x}_{n}]^{\rm T}|{x}_{i}\in\{\underline{x}_{i},\overline{x}_{i}\}\}$. In this work, $n$-dimensional real spaces are denoted as $\mathbb{R}^{n}$, $\mathbb{N}$ is used to represent the set of natural numbers, and $\mathbb{N}_{n}:=\{1,2,\ldots,n\}$. Let $[a,b]:=\{a,\ldots,b\}$ for $a,b\in\mathbb{N}$ be a shorthand notation for a continuous set of natural numbers. Similarly, let ${\bm{x}}_{[a,b]}:=\{x_{a},\ldots,x_{b}\}$ be a shorthand notation for a set consisting of items of the vector ${\bm{x}}=[x_{a},\ldots,x_{b}]^{\rm T}$. The $\mathcal{L}_{2}$-norm of a function is denoted as $\|\cdot\|_{\mathcal{L}_{2}}$, and the $\mathcal{L}_{2}$-norm of a vector-valued function ${\bm{g}}=[g_{1},\ldots,g_{m}]^{\rm T}$ is defined as $\|{\bm{g}}\|_{\mathcal{L}_{2}}^{2}:=\sum\limits_{i=1}^{m}\|g_{i}\|^{2}_{\mathcal{L}_{2}}$.
II Problem Formulation
Consider a nonlinear dynamic system of the form
${\bm{x}}(k+1)={\bm{f}}({\bm{x}}(k),{\bm{u}}(k),{\bm{v}}(k)),$ (1)
where ${\bm{x}}(k)\in\mathcal{X}\subset\mathbb{R}^{n}$ represents the state, ${\bm{v}}(k)$ denotes unknown disturbances, ${\bm{u}}(k)\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, and ${\bm{f}}(\cdot)=[f_{0}(\cdot),\ldots,f_{n}(\cdot)]^{\rm T}$ are unknown system dynamics. In this work, the control input is generated by a controller ${\bm{g}}(\cdot)$, i.e., ${\bm{u}}(k)={\bm{g}}({\bm{x}}(k))=[g_{1}({\bm{x}}(k)),\ldots,g_{m}({\bm{x}}(k))]^{\rm T}\in\mathbb{R}^{m}$.
The control objective of this work is to design a controller ${\bm{g}}(\cdot)$ for the given unknown system (1) and a specific task, such that a cost function $J$ is minimized. In practice, the cost can be computed based on the system states and inputs over time instants $\{0,\ldots,K\}$. A widely used example is the quadratic cost function $J=\sum\limits_{k=0}^{K}({\bm{x}}(k)^{\rm T}{Q}{\bm{x}}(k)+{\bm{u}}(k)^{\rm T}{R}{\bm{u}}(k))$, where ${Q}$ and ${R}$ are positive definite matrices. Another commonly discussed example is the exponential cost function (also referred to as the risk-sensitive cost function [34]) $J=\sum\limits_{k=0}^{K}e^{-\alpha k}({\bm{x}}(k)^{\rm T}{Q}{\bm{x}}(k)+{\bm{u}}(k)^{\rm T}{R}{\bm{u}}(k))$, where $\alpha>0$ is a user-defined parameter. In this work, we assume that the cost function can also be computed from the system states and input data, as illustrated in the aforementioned examples.
For the sake of analysis, we provide an alternative perspective to describe the cost function $J$. For a given system and control task, the cost function $J$ can be regarded as a random function with respect to the controller ${\bm{g}}(\cdot)$. Therefore, with a slight abuse of notation, the cost is expressed as $J({\bm{g}})$. The controller design problem can then be formulated as solving the optimization problem $\min\limits_{{\bm{g}}\in\mathcal{G}}J({\bm{g}})$, where $\mathcal{G}$ is the set of all possible controllers.
Considering that searching for a control law in the infinite function space $\mathcal{G}$ is an impractical task, especially when the system dynamics $f(\cdot)$ are unavailable, this paper proposes a learning method based on an initial control law. This is feasible in engineering practice, as many devices normally operate using control laws obtained through empirical or other design methods.
For an initial control law $\check{\bm{g}}$, we aim to extract its structural information to construct a Hilbert space $\check{\mathcal{G}}$ such that $\check{\bm{g}}\in\check{\mathcal{G}}$. The Hilbert space is chosen because it has good properties that facilitate the exploration of potential control laws to improve the system’s control performance. For example, thanks to the inner product structure of the Hilbert space, orthogonal function bases can be used to represent possible control laws; the distance metric in the Hilbert space allows us to discuss the proximity of different control laws, enabling methods such as gradient descent to solve optimization problems. For a cost function $J$, online data is collected to explore potential optimal control laws to minimize $J$, i.e., $\min\limits_{{\bm{g}}\in\check{\mathcal{G}}}J$. To facilitate the discussion, assumptions about the initial control law and the cost function are introduced as follows:
Assumption 1.
There exists a set $\bar{\mathcal{X}}$ satisfying:
  1. 1.
    $\mathcal{X}\subseteq\bar{\mathcal{X}}$;
  2. 2.
    $\forall~{}\overline{\bm{x}},\underline{\bm{x}}\in\bar{\mathcal{X}}:\overline{\bm{x}}\otimes\underline{\bm{x}}\subset\bar{\mathcal{X}}$.
Assumption 2.
The cost function $J$ is Lipschitz continuous with Lipschitz constant $L_{J}$, i.e.,
$|J({\bm{g}}_{1})-J({\bm{g}}_{2})|\leq L_{J}\|{\bm{g}}_{1}-{\bm{g}}_{2}\|_{\mathcal{L}_{2}},{\bm{g}}_{1},{\bm{g}}_{1}\in\check{\mathcal{G}},$ (2)
and $J({\bm{g}})$ is bounded as $J({\bm{g}})\geq\underline{J_{m}},{\bm{g}}\in\check{\mathcal{G}}$.
Remark 1.
Assumption 1 can be easily satisfied by constructing an appropriate $\bar{\mathcal{X}}$. For example, for any set $\mathcal{X}$, the set $\bar{\mathcal{X}}$ can be defined as: $\bar{\mathcal{X}}=\{{\bm{x}}|\|{\bm{x}}\|_{\infty}\leq\max\limits_{\bar{\bm{x}}\in\mathcal{X}}\|\bar{\bm{x}}\|_{\infty}\}$. The purpose of Assumption 1 is to simplify the structure of the set $\mathcal{X}$ to facilitate the design of the control law. Assumption 2 can also be satisfied when considering bounded sets $\mathcal{X},\mathcal{U}$, which aligns with practical scenarios because of the bounded energy restrictions in practical systems.
To avoid unnecessary computational burden or potential system oscillations caused by frequent updates, a segment-based learning method is adopted. Concretely, system data is divided into different segments, and the data of a segment is used as a whole for learning. The control law during system operation is updated at the beginning of each segment. Specifically, let $K$ be the segment length, and $t$ be used to represent the segment index, then the dynamic characteristics of the system can be expressed as:
${\bm{x}}_{t}(k)$ $=f({\bm{x}}_{t}(k-1),{\bm{u}}_{t}(k-1),{\bm{v}}_{t}(k)),$ $0\leq k\leq K.$ (3)
The controller used in the $t$-th segment is written as ${\bm{g}}_{t}$.
With the aforementioned scenario, this work discusses the following three questions:
  1. 1.
    How to construct and evaluate the function space $\check{\mathcal{G}}$?
  2. 2.
    How to update the control law ${\bm{g}}_{t}$ to balance exploration and exploitation?
  3. 3.
    How to evaluate the closed-loop performance of the learning-based control systems?
III Preliminaries
In this section, some useful notations and definitions are introduced for use in subsequent analysis.
Definition 1.
(Hellinger distance [35]) Let $p_{1}$ and $p_{2}$ be probability densities on a measurable space $(\mathcal{P},\Sigma_{p})$ w.r.t. a measure $v$, then the Hellinger distance between $p_{1}$ and $p_{2}$ is defined as
$D_{\rm H}(p_{1},p_{2})=\frac{1}{\sqrt{2}}\left[\int_{\mathcal{P}}\left(\sqrt{p_{1}(w)}-\sqrt{p_{2}(w)}\right)^{2}{\rm d}v(w)\right]^{\frac{1}{2}}.$ (4)
Definition 2.
(Kullback-Leibler divergence, KL divergence, [35]) Let $p_{1}$ and $p_{2}$ be probability densities on a measurable space $(\mathcal{P},\Sigma_{p})$ w.r.t. a measure $v$, then the Kullback-Leibler divergence from $p_{1}$ to $p_{2}$ is defined as
$D_{\rm KL}(p_{1}||p_{2})=\int_{\mathcal{P}}p_{1}(w)\log\left(\frac{p_{1}(w)}{p_{2}(w)}\right){\rm d}v(w).$ (5)
Definition 3.
(Neighborhood [31]) Consider a probability density function $p$ on the measurable space $(\mathcal{P},\Sigma_{p})$ with metric $d$, a neighborhood with center $p_{1}$ and radius $\delta$ is defined as
$\mathcal{B}(p_{1},\delta):=\{p\in\mathcal{P}:d(p,p_{1})<\delta\},$ (6)
which complementary set is defined as
$\mathcal{B}^{\rm c}(p_{1},\delta):=\{p\in\mathcal{P}:d(p,p_{1})\geq\delta\}.$ (7)
Remark 2.
Neighborhoods are used to represent a small region around an element. They consist of three key components: a given element (center) $p_{1}$, a metric (distance) $d$, and a radius (size) $\delta$. In this work, unless otherwise specified, the symbol $\mathcal{B}$ represents neighborhoods under either the Hellinger distance or the KL divergence. When a specific metric is indicated, $\mathcal{B}_{\rm H}(\cdot,\cdot)$ denotes neighborhoods under the Hellinger distance, and $\mathcal{B}_{\rm KL}(\cdot,\cdot)$ denotes neighborhoods under the KL divergence. A potential caveat here is that the KL divergence is not a true metric, as it is asymmetric (i.e., $D_{\rm KL}(p_{1}||p_{2})\neq D_{\rm KL}(p_{2}||p_{1})$) and does not satisfy the triangle inequality. However, when the center $p_{1}$ is specified, Definition 3 is still meaningful for the KL divergence, and thus it is also discussed in this work.
IV Main Results
The proposed learning approach is introduced in this section, which is divided into four parts. The first part proposes a parameterization method for the space $\mathcal{G}$ and discusses the relationship between the constructed function space and the optimal control law. Then, a data-driven method is designed to explore the optimal control law within the function space. The third subsection discusses the convergence of the learning process. The last part analyzes the closed-loop performance of the system during the learning process, characterizing the expected system cost through the regret compared to the optimal control law.
IV-A Parameterization of Controllers
In this work, the control law ${\bm{g}}(\cdot)$ is chosen from a function space $\{g|g:\mathcal{X}\rightarrow\mathcal{U}\}$. Before introducing the function space construction method proposed in this work, we would like to point out that there are many potential feasible construction methods. When partial information about the system dynamics is available, the corresponding function space $\check{\mathcal{G}}$ can be naturally obtained. However, in this work, we focus on the case where the system dynamics are unknown. In this case, the function space $\check{\mathcal{G}}$ is constructed using the initial control law $\check{\bm{g}}(\cdot)$. Specifically, Hilbert spaces $\check{\mathcal{G}}_{i},i\in\{1,\ldots m\}$ are constructed based on the initial control law, and convex hulls $\mathcal{G}_{i}\subset\check{\mathcal{G}}_{i}$ are used to learn the control law. The proposed method is inspired by the fact that many existing control laws are designed based on empirical or other design methods, which can be used as a starting point for constructing the function space.
For a vector $\tilde{\bm{x}}\in\mathcal{X}$ and a set ${\bm{w}}\subseteq[1,n]$, we denote a vector ${\bm{d}}=[d_{1},\ldots,d_{n}]=:({\bm{x}},\tilde{\bm{x}})_{\bm{w}}$ as follows:
$d_{i}=\left\{\begin{array}[]{ll}x_{i},{}&{\text{if}}~{}i\in{\bm{w}},\\ \tilde{x}_{i},{}&{\text{if}}~{}i\notin{\bm{w}}.\end{array}\right.$ (10)
Then, a proposition on $({\bm{x}},\tilde{\bm{x}})_{\bm{w}}$ is proposed as follows.
Proposition 1.
If Assumption 1 holds, $\forall{\bm{x}},\tilde{\bm{x}}\in\mathcal{X}$ and $\forall{\bm{w}}\subseteq[1,n]$, $({\bm{x}},\tilde{\bm{x}})_{\bm{w}}\in\bar{\mathcal{X}}$.
Proof.
The proof can be achieved by utilizing Assumption 1 and the definition of $\bar{\mathcal{X}}$, and is therefore omitted here.
Using the notations aforementioned and a given initial control law $\bar{\bm{g}}(\cdot)=[\bar{g}_{1}(\cdot),\ldots,\bar{g}_{m}(\cdot)]^{\rm T}$, several sets $\mathcal{G}_{1},\ldots,\mathcal{G}_{m}$ are first introduced as
$\mathcal{G}_{i}:=\Bigg{\{}{g}\Bigg{|}$ ${g}=\sum\limits_{{\bm{w}}\subseteq[1,n]}\alpha({\bm{w}})\Gamma g^{(i)}_{{\bm{w}},\tilde{\bm{x}}},\alpha({\bm{w}})\geq 0,$
$\sum\limits_{{\bm{w}}\subseteq[1,n]}\alpha({\bm{w}})=1,\Bigg{\}},~{}i\in\{1,\ldots,m\},$ (11)
where $\Gamma>0$ is a user-defined constant, and $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}$ is a function that depends on the variables listed in ${\bm{w}}$ and a constant $\tilde{\bm{x}}$, which satisfies
  1. i)
    $g^{(i)}_{\emptyset,\tilde{\bm{x}}}=\check{g}_{i}(\tilde{\bm{x}}),~{}g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}({\bm{x}})=\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{\bm{w}})-\sum\limits_{{\bm{\varpi}}\subsetneq{\bm{w}}}g^{(i)}_{{\bm{\varpi}},\tilde{\bm{x}}}({\bm{x}})$;
  2. ii)
    $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}({\bm{x}})=\sum\limits_{\varpi\subseteq w}(-1)^{|{\bm{w}}|-|{\bm{\varpi}}|}g_{i}(({\bm{x}},\tilde{\bm{x}})_{\bm{\varpi}})$;
  3. iii)
    $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}({\bm{x}})=0,\text{if}~{}{\bm{x}}=\tilde{\bm{x}}$,
with $|{\bm{w}}|$ representing the number of elements in ${\bm{w}}$. The set $\mathcal{G}_{i}$ can be regarded as the convex hull of multiple points in the function space $\check{\mathcal{G}}_{i}$ composed of basis functions $\{g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}\}$.
The parameterization algorithm for the set $\mathcal{G}_{i}$ is summarized in Algorithm 1. The algorithm is organized as a recursive function that computes the basis functions $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}$ based on the initial control law $\check{g}_{i}(\cdot)$ and the constant $\tilde{\bm{x}}$. The algorithm starts with the empty set ${\bm{w}}=\emptyset$, where it directly returns the value of the initial control law at $\tilde{\bm{x}}$. For other cases, it recursively computes the basis functions by summing over all proper subsets of ${\bm{w}}$, ensuring that the contribution of different combinations of variables can be captured by the constructed basis functions.
Algorithm 1 Parameterization Algorithm for $\mathcal{G}_{i}$.
1:  Function: Output=Func($\check{g}_{i}$,${\bm{w}}$,$\tilde{\bm{x}}$);
2:  if ${\bm{w}}=\emptyset$ then
3:     $g^{(i)}_{\emptyset,\tilde{\bm{x}}}=\check{g}_{i}(\tilde{\bm{x}})$;
4:     Output=$g^{(i)}_{\emptyset,\tilde{\bm{x}}}$;
5:  else
6:     $g^{(i)}=0$;
7:     for each ${\bm{\varpi}}\subsetneq{\bm{w}}$ do
8:        $g^{(i)}=g^{(i)}$+Func($\check{g}_{i}$,${\bm{\varpi}}$,$\tilde{\bm{x}}$)
9:     end for
10:     Output=$\check{g}_{i}(({\bm{x}},\tilde{\bm{x}})_{\bm{w}})-g^{(i)}$;
11:  end if
Furthermore, the set $\mathcal{G}$ is defined as the Cartesian product of sets $\mathcal{G}_{i},~{}i\in\{1,\ldots,m\}$ as follows:
$\mathcal{G}$ $:=\mathcal{G}_{1}\times\mathcal{G}_{2}\times\ldots\times\mathcal{G}_{m}.$ (12)
In this work, we assum that the $\mathcal{L}_{2}$-norm of the considered base functions are bounded, i.e., $\mathcal{G}_{i}\subset\mathcal{G}^{*}$ with $\mathcal{G}^{*}$ defined as
$\mathcal{G}^{*}=\left\{g|\|g\|_{\mathcal{L}_{2}}\leq M_{g}\right\}.$ (13)
To analyze the learning error, the following assumption for the optimal controller is adopted.
Assumption 3.
For the considered system (1), there exists an optimal controller ${\bm{g}}^{*}=[g_{1}^{*},\ldots,g_{m}^{*}]^{\rm T},~{}g_{i}^{*}\in\mathcal{G}^{*}$ that minimizes the cost function $J({\bm{g}})$.
Remark 3.
Assumption 3 is used to ensure the existence of the optimal control law. On one hand, in the field of signal processing and control theory, the integral of the square of a signal is often regarded as the energy of the signal. In practice, the bounded energy of most systems makes it natural to consider the $\mathcal{L}_{2}$ space. Therefore, Assumption 3 only adds a weak premise for the selection of the control law ${\bm{g}}$. On the other hand, the fact that the square-integrable space is a Hilbert space provides structural advantages for discussing the properties of the optimal solution.
In the following, we show that the constructed function space can be quantitatively evaluated by analyzing the regret of the optimal control law $\bar{\bm{g}}=\arg\min\limits_{{\bm{g}}\in\mathcal{G}}J$ with respect to to the optimal control law ${\bm{g}}^{*}$ in Assumption 3.
Lemma 1.
(Lemma 1 in [36]) If ${\bm{g}}^{*}$ is in the closure of the convex hull of a set $G$ in a Hilbert space, with $\|{\bm{g}}\|_{\mathcal{L}_{2}}\leq M$ for all ${\bm{g}}\in G$, there is a $\bar{\bm{g}}_{n}$ in the convex hull of $n$ points in $G$ such that
$\|{\bm{g}}^{*}-\bar{\bm{g}}_{n}\|_{\mathcal{L}_{2}}^{2}\leq\frac{M^{2}-\|{\bm{g}}^{*}\|_{\mathcal{L}_{2}}}{n}$ (14)
Definition 4.
A reproducing kernel Hilbert space (RKHS) $\mathcal{H}(K)$ is a Hilbert space of functions on a set $\mathcal{X}$ with a reproducing kernel $K(\cdot,\cdot)$ that satisfies the following properties:
  1. i)
    $K({\bm{x}},\cdot)\in\mathcal{H}(K)$ for all ${\bm{x}}\in\mathcal{X}$;
  2. ii)
    $K({\bm{x}}_{i},{\bm{x}}_{j})=K({\bm{x}}_{j},{\bm{x}}_{i})$ for all ${\bm{x}}_{i},{\bm{x}}_{j}\in\mathcal{X}$;
  3. iii)
    $\langle g,K({\bm{x}},\cdot)\rangle_{\mathcal{H}(K)}=g({\bm{x}})$ for all ${\bm{x}}\in\mathcal{X}$ and $g\in\mathcal{H}(K)$,
[…]
With the above tools, we are ready to present our first main result.
Theorem 1.
Let Assumptions 1 and 3 hold. Under the proposed parameterization approach (12), the constrcted learning set $\mathcal{G}$ satisifes: $\exists\bar{\bm{g}}=\arg\min\limits_{{\bm{g}}\in\mathcal{G}}J$:
$\|\bar{g}_{i}-{g}_{i}^{*}\|_{\mathcal{L}_{2}}\leq\sqrt{\frac{M_{g}^{2}-\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}{2^{n}}},~{}\forall i\in\{1,\ldots,m\}.$ (15)
Proof.
The proof of Theorem 1 can be easily completed by analyzing the properties of the set $\mathcal{G}_{i}$.
Specifically, according to the definition of $\mathcal{G}_{i}$ in 4, it can be verified that for $\forall{\bm{w}}\subseteq[1,n]$, the function $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}$ is in the RKHS $\mathcal{H}(K_{n,{\bm{w}}}^{(i)})$ with the reproducing kernel $K_{n,{\bm{w}}}^{(i)}({\bm{x}},\tilde{\bm{x}})$. Here, $\mathcal{H}(K_{n,{\bm{w}}}^{(i)})$ is a Hilbert space of real functions defined on $\mathcal{X}_{\bm{w}}:=\{{\bm{x}}_{\bm{w}}|x\in\mathcal{X}\}$.
According to the definition (IV-A), functions $g\in\mathcal{G}_{i}$ can be regarded as linear combinations of base functions $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}},~{}{\bm{w}}\subseteq[1,n]$. Therefore, for $g\in\mathcal{G}_{i}$, a RKHS can be obtained as $\mathcal{H}(K_{n}^{(i)})$ with the reproducing kernel $K_{n}^{(i)}({\bm{x}},\tilde{\bm{x}})$ being
$K_{d}^{(i)}({\bm{x}},\tilde{\bm{x}})=\sum\limits_{{\bm{w}}\subseteq[1,n]}K_{n,{\bm{w}}}^{(i)}({\bm{x}}_{\bm{w}},\tilde{\bm{x}}_{\bm{w}}),K_{n,\emptyset}\equiv 1.$ (16)
As a result, the set $\mathcal{G}_{i}$ can be expressed as the convex hull of $2^{n}$ points in the RKHS $\mathcal{H}(K_{n}^{(i)})$. Therefore, according to Lemma 1, there exists a $\bar{\bm{g}}_{i}$ in the constructed set $\mathcal{G}_{i}$ satisfying
$\|\bar{g}_{i}-{g}_{i}^{*}\|^{2}_{\mathcal{L}_{2}}\leq{\frac{M_{g}^{2}-\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}{2^{n}}},~{}\forall i\in\{1,\ldots,m\},$ (17)
which completes the proof of Theorem 1.
Remark 4.
Theorem 1 analyzes the regret introduced by the proposed controller parameterization method. Assumption 3 assumes an optimal control law ${\bm{g}}^{*}$ in the space $\mathcal{G}^{*}$. Meanwhile, the parameterization method (12) constructs a convex hull $\mathcal{G}_{i}$. The reduction of the feasible space inevitably leads to changes in the optimal solution of the optimization problem, which is quantified in Theorem 1 through the regret $\|{\bm{g}}^{*}-\bar{\bm{g}}_{n}\|_{\mathcal{L}_{2}}$. Theorem 1 demonstrates that the proposed controller parameterization method ensures the boundedness of the regret.
Remark 5.
The regret of the parameterization method is guaranteed by the properties of the Hilbert space (Lemma 1). Therefore, the proof of Theorem 1 primarily aims to demonstrate that the constructed function space is a Hilbert space. The proposed approach not only ensures the boundedness of the learning regret through Lemma 1 but also provides an advantageous structural foundation for solving optimization problems within the function space.
Using Theorem 1, the error bound for vector-valued functions can be further derived, which is presented in the following corollary.
Corollary 1.
Let Assumptions 1 and 3 hold. Under the proposed parameterization approach (12), the constructed learning set $\mathcal{G}$ satisfies: $\exists\bar{\bm{g}}=\arg\min\limits_{{\bm{g}}\in\mathcal{G}}J$:
$\|\bar{\bm{g}}-{\bm{g}}^{*}\|^{2}_{\mathcal{L}_{2}}\leq\frac{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}{2^{n}}.$ (18)
Proof.
Corollary 1 can be directly derived from Theorem 1 and the definition of $\mathcal{L}_{2}$-norm of vector-valued functions. Thus the proof is omitted here.
IV-B TS-based Learning Algorithm
This subsection introduces an online active sampling method for control laws. Specifically, the proposed method is a policy-oriented ALC approach. Since ${\bm{g}}(\cdot)$ is not known, evaluating the performance of control laws is challenging. Therefore, this work first designs a TS method to learn the function $J({\bm{g}})$. At the beginning of each segment, the control law ${\bm{u}}_{t}(k)={\bm{g}}_{t}({\bm{x}}_{t}(k))$ is selected as the one that minimizes the currently learned function, which is denoted as $J_{t}({\bm{g}})$.
In the following, we first introduce the space used to learn the function $J({\bm{g}})$. Then, an online update method is proposed within a Bayesian framework. To further explain the learning process, pseudocode and corresponding explanations are also provided in this subsection.
For the considered system (1) and a given data segment $\{{\bm{x}}_{t}(0),{\bm{u}}_{t}(0),\ldots,{\bm{x}}_{t}(K),{\bm{u}}_{t}(K)\}$, the control cost can be calculated according to its definition. However, due to the unknown system dynamics $f(\cdot)$ and uncertain noise ${\bm{w}}$, the cost function $J$ is not available unless the control law is executed. This brings challenges to the controller design. To address this issue, we consider a function space $\bar{\mathcal{J}}$ with $\sigma$-field $\Sigma_{\bar{\mathcal{J}}}$ and measure $v_{\bar{\mathcal{J}}}(\cdot)$ such that the function $\bar{J}({\bm{g}};{\bm{x}})\in\bar{\mathcal{J}}$ satisfies the following properties:
  1. 1.
    Non-negativity:
    $\bar{J}({\bm{g}};{\bm{x}})>0,~{}{\bm{g}}\in\mathcal{G},~{}{\bm{x}}\in\mathcal{X};$ (19)
  2. 2.
    Normalization:
    $\int_{\mathcal{G}}\bar{J}({\bm{g}};{\bm{x}}){\rm d}v_{\bar{\mathcal{J}}}({\bm{g}})=1;$ (20)
  3. 3.
    Probability property:
    $\mathbb{P}({\bm{g}}\in\Omega)=\int_{\Omega}\bar{J}({\bm{g}};{\bm{x}}){\rm d}v_{\bar{\mathcal{J}}}({\bm{g}}),~{}\forall\Omega\subseteq\mathcal{G};$ (21)
  4. 4.
    Proportion:
    $\bar{J}({\bm{g}};{\bm{x}})\propto\frac{1}{J},~{}{\bm{g}}\in\mathcal{G},{\bm{u}}(k)={\bm{g}}({\bm{x}}(k)),{\bm{x}}(0)={\bm{x}}.$ (22)
  5. 5.
    Boundedness: $v_{\bar{\mathcal{J}}}(\bar{\mathcal{J}})=\bar{M}_{\bar{\mathcal{J}}},~{}\underline{J_{b}}\leq\bar{J}({\bm{g}};{\bm{x}})\leq\overline{J_{b}}.$
Here we note that, through the aforementioned five properties, we establish a space $\bar{\mathcal{J}}$, which serves as a tool for indirectly learning the unknown cost function $J(\cdot)$. Although there are various possibilities for the type of the unknown cost function, under the boundedness assumption 2, we can construct a function $\bar{J}({\bm{g}};{\bm{x}})$ belonging to the space $\bar{\mathcal{J}}$ and related to the cost function $J(\cdot)$. This facilitates the following learning mechanism design and learning convergence analysis.
Through the above construction, we obtain a function that characterizes the distribution of the control law ${\bm{g}}$. Due to property (22), the control law ${\bm{g}}$ is more likely to be distributed in the region where the cost function $J$ is relatively small. In fact, for a given dynamic characteristic $f(\cdot)$, the distribution of the cost function $J$ can be determined by the control law ${\bm{g}}$ and the initial state of the system ${\bm{x}}(0)$. We assume the initial state is consistent across segments, i.e., ${\bm{x}}_{t_{i}}(0)={\bm{x}}_{t_{j}}(0)$. This can be achieved by designing a reference signal or adjusting segment start times to begin when the system state reaches a predefined value. Thus, when it does not cause confusion, $\bar{J}({\bm{g}};{\bm{x}})$ will be abbreviated as $\bar{J}({\bm{g}})$.
In this work, we will indirectly learn the cost function $J({\bm{g}})$ by learning the constructed function $\bar{J}({\bm{g}})$. Specifically, let $F_{\rm pdf}(\bar{J})$ be the probability density function (pdf) of the function $\bar{J}$ in the space $\bar{\mathcal{J}}$. The prior distribution is denoted as $F^{0}_{\rm pdf}(\bar{J})$, and can be updated by the following formula:
$F_{\rm pdf}^{t}(\bar{J})=\frac{R_{t}F_{\rm pdf}^{0}(\bar{J})}{\int_{\bar{\mathcal{J}}}R_{t}F_{\rm pdf}^{0}({\rm d}v_{\mathcal{J}}(\bar{J}))}$ (23)
where $R_{t}$ is defined as
$R_{t}:=\prod\limits_{i=1}^{t}\frac{\bar{J}({\bm{g}}_{t})}{\bar{J}^{*}({\bm{g}}_{t})},$ (24)
with $\bar{J}^{*}({\bm{g}})$ being
$\bar{J}^{*}({\bm{g}})\propto\frac{1}{J_{t}({\bm{g}})}.$ (25)
In the above equation, $J_{t}({\bm{g}})$ is the observed cost of the system at $t$-th segment with the control law being ${\bm{g}}$.
For a subset $\Omega\subseteq\bar{\mathcal{J}}$ and a time instant $t$, the probability $\mathbb{P}(\bar{J}\in\Omega)$ can be calculated using $F_{\rm pdf}^{t}(\bar{J})$ as
$\mathbb{P}(\bar{J}\in\Omega)=\int_{\Omega}F_{\rm pdf}^{t}(\bar{J}){\rm d}v_{\mathcal{J}}(\bar{J})=\frac{\int_{\Omega}R_{t}F_{\rm pdf}^{0}({\rm d}v_{\mathcal{J}}(\bar{J}))}{\int_{\bar{\mathcal{J}}}R_{t}F_{\rm pdf}^{0}({\rm d}v_{\mathcal{J}}(\bar{J}))}.$ (26)
Based on the above equation, the predictive density at the $t$-th segment can be obtained as the expectation defined as
$\bar{J}^{t}:=\mathbb{E}\left[\bar{J}\right]:=\int_{\bar{\mathcal{J}}}\bar{J}F_{\rm pdf}^{t}({\rm d}v_{\mathcal{J}}(\bar{J})).$ (27)
Algorithm 2 TS-based ALC.
1:  Input: prior distribution $F_{\rm pdf}^{0}(\bar{J})$;
2:  Randomly sample a function $\bar{J}_{0}$ with initial pdf $F_{\rm pdf}^{0}(\bar{J})$;
3:  Select an initial control law by solving optimization problem ${\bm{g}}_{0}=\arg\max\limits_{{\bm{g}}\in\mathcal{G}}\bar{J}_{0}({\bm{g}})$;
4:  Apply the control law ${\bm{g}}_{0}$ to the system and observe sample data in the initial segment
5:  Initialize $R_{0}=1$;
6:  for each iteration $t=1,2,\ldots$ do
7:     Update the posterior distribution:
$R_{t}$ $=\frac{\bar{J}({\bm{g}}_{t-1})}{\bar{J}^{*}({\bm{g}}_{t-1})}R_{t-1};$
$F_{\rm pdf}^{t}(\bar{J})$ $=\frac{R_{t}F_{\rm pdf}^{0}(\bar{J})}{\int_{\bar{\mathcal{J}}}R_{t}F_{\rm pdf}^{0}({\rm d}v_{\mathcal{J}}(\bar{J}))};$
8:     Randomly sample $\bar{J}_{t}$ according to posterior distribution $F_{\rm pdf}^{t}(\bar{J})$;
9:     Select the controller ${\bm{g}}_{t}$ that maximizes $\bar{J}_{t}$;
10:     Apply the controller ${\bm{g}}_{t}$ and collect data in the $t$-th segment;
11:  end for
Using the proposed distribution function $F_{\rm pdf}^{t}(\bar{J})$, the TS based learning algorithm can be designed and is summarized in Algorithm 2. After obtaining the initial data (Lines 1-5), the control law is updated at the beginning of each segment. Specifically, at the beginning of each segment, the control cost $J_{t-1}({\bm{g}}_{t-1})$ is calculated using the data collected in the previous segment, and $R_{t}$ and the posterior distribution function $F_{\rm pdf}^{t}(\bar{J})$ are updated (Line 7). Here, a small trick is employed: although $\bar{J}^{*}({\bm{g}}_{t})$ and $1/J_{t}({\bm{g}}_{t})$ differ by a proportional coefficient (as shown in (25)), this coefficient appears in both the numerator and denominator when updating the posterior distribution, so it does not affect the calculation result. Then, a function $\bar{J}_{t}$ is randomly sampled according to the posterior distribution function $F_{\rm pdf}^{t}(\bar{J})$ (Line 8), and the control law ${\bm{g}}_{t}$ is selected to maximize $\bar{J}_{t}({\bm{g}}_{t})$ (Line 9). Finally, the control law ${\bm{g}}_{t}$ is applied, and the system data for the $t$-th segment $\{{\bm{x}}_{t}(0),{\bm{u}}_{t}(0),\ldots,{\bm{x}}_{t}(K),{\bm{u}}_{t}(K)\}$ is collected (Line 10).
The proposed learning method is an end-to-end control strategy that directly learns the control laws from sampled data, without requiring system model identification and parameter estimation. According to system identification theories [37], system identification typically assumes that the true system dynamics can be described by a certain class of models. This inevitably limits the design of control laws, as the design is tailored to the model rather than the system itself. The proposed method directly learns the relationship between control laws and costs through random sampling, avoiding assumptions and restrictions on the system model. Thanks to its flexibility, the proposed method can be applied to various types of systems without making assumptions about the system and the distribution of the noise ${\bm{v}}$.
IV-C Convergence Analysis of the Learning Process
This subsection discusses the relationship between the function $\bar{J}_{t}(\cdot)$ obtained under the proposed learning method and the true reward $\bar{J}^{*}(\cdot)$. To aid our analysis, we write
${L_{t}(\Omega)}={\int_{\Omega}R_{t}F_{\rm pdf}^{0}({\rm d}v_{\mathcal{J}}(\bar{J}))}$ (28)
for convenience. For KL divergence $D_{KL}$ and Hellinger distance $D_{H}$, we define a function $T_{d}(x)$ as
$T_{d}(x)$ :=$:=\left\{\!\!\begin{array}[]{ll}\log x,&\text{if}~{}d=D_{\rm KL},\\ \sqrt{x}-1,&\text{if}~{}d=\frac{1}{2}D^{2}_{\rm H}.\end{array}\right.$ (31)
Furthermore, by combining equations (28) and (27), we can define a function $T_{d}^{t}$ as
$T_{d}^{t}:=T_{d}\left(\frac{L_{t}({\mathcal{B}}^{c}(\bar{J}^{*},\delta_{\bar{J}}))}{L_{t-1}({\mathcal{B}}^{c}(\bar{J}^{*},\delta_{\bar{J}}))}\right),$ (32)
where ${\mathcal{B}}(\bar{J}^{*},\delta_{\bar{J}})$ is a neighborhood of the predictive density $\bar{J}^{*}$ with radius $\delta_{\bar{J}}$, which can be defined on eigher KL divergence $d=D_{KL}$ or Hellinger distance $d=\frac{1}{2}D^{2}_{\rm H}$. Then, the non-asymptotic convergence property of the predictive density can be summarized in the following theorem.
Theorem 2.
The predictive density $\bar{J}^{T}$ obtained in (27) satisfies
[…]
if the following two conditions are satisfied:
$\lim\limits_{t\rightarrow\infty}\inf d(\bar{J}^{t},\bar{J}^{*})>0;$ (34)
$\sum\limits_{t=1}^{T}t^{-2}\mathbb{V}(T_{d}^{t})<\infty.$ (35)
Proof.
The proof of this theorem is divided into two steps. In the first step, a martingale process is constructed, and several properties of the martingle are analyzed. In the second step, the proof of Theorem 2 is completed by using the properties of the constructed martingale process. In the following analysis, the complementary set of the neighborhood ${\mathcal{B}}(\bar{J}^{*},\delta_{\bar{J}})$ is denoted as $\Omega={\mathcal{B}}^{c}(\bar{J}^{*},\delta_{\bar{J}})$.
Step1: martingale analysis.  We consider the $\sigma$-field
$\Sigma_{\bar{{J}}}^{t}=\sigma(\bar{J}_{1},\ldots,\bar{J}_{t})=\sigma(\bar{J}({\bm{g}}_{1}),\ldots,\bar{J}({\bm{g}}_{t})),$ (36)
which is generated by the control costs of segments $1,\ldots,t$.
According to the definition of $L_{t}$, we have
$\frac{L_{t+1}(\Omega)}{L_{t}(\Omega)}=\frac{\int_{\Omega}R_{t+1}F_{\rm pdf}^{0}({\rm d}{v_{\mathcal{J}}(\bar{J})})}{\int_{\Omega}R_{t}F_{\rm pdf}^{0}({\rm d}{v_{\mathcal{J}}(\bar{J})})}$ (37)
$=$ $\frac{\int_{\Omega}\bar{J}({\bm{g}}_{t+1})R_{t+1}F_{\rm pdf}^{0}({\rm d}{v_{\mathcal{J}}(\bar{J})})}{\int_{\Omega}\bar{J}({\bm{g}}_{t+1})R_{t}F_{\rm pdf}^{0}({\rm d}{v_{\mathcal{J}}(\bar{J})})}=\frac{\bar{J}^{t}({\bm{g}}_{t+1})}{\bar{J}^{*}({\bm{g}}_{t+1})}.$ (38)
Then, for the considered $\sigma$-field $\Sigma_{\bar{{J}}}^{t}$, the following equations can be verified for Hellinger distance $D_{\rm H}(\cdot,\cdot)$:
$\mathbb{E}\left[\left.\sqrt{\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}}-1\right|\Sigma_{\bar{J}}^{t}\right]=\mathbb{E}\left[\left.\sqrt{\frac{\bar{J}^{t}({\bm{g}}_{t})}{\bar{J}^{*}({\bm{g}}_{t})}}-1\right|\Sigma_{\bar{J}}^{t}\right]$
$=$ $-\frac{1}{2}D_{\rm H}^{2}(\bar{J}^{t}(\cdot),\bar{J^{*}}(\cdot)).$ (39)
Similarly, for KL divergence $D_{\rm KL}(\cdot,\cdot)$, we have
$\mathbb{E}\left[\left.\log\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}\right|\Sigma_{\bar{J}}^{t}\right]=\mathbb{E}\left[\left.\log\frac{\bar{J}^{t}({\bm{g}}_{t})}{\bar{J}^{*}({\bm{g}}_{t})}\right|\Sigma_{\bar{J}}^{t}\right]$
$=$ $-D_{\rm KL}(\bar{J}^{t}_{\Omega}(\cdot)||\bar{J^{*}}(\cdot)).$ (40)
By recalling the definition of $T_{d}(\cdot)$ in (31), the following equation can be verified for both Hellinger distance and KL divergence:
$\mathbb{E}\left[\left.T_{d}\left(\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}\right)\right|\Sigma_{\bar{{J}}}^{t}\right]=-d(\bar{J}^{t}({\bm{g}}),\bar{J}^{*}({\bm{g}})).$ (41)
Then according to the definition of $T_{d}^{t}$ defined in (32), equation (41) can be equivalently rewritten as the following form for both Hellinger distance and KL divergence:
$\mathbb{E}[T_{d}^{t}]+d(\bar{J}^{t}(\cdot),\bar{J}^{*}(\cdot))=0.$ (42)
For the stochastic process $M_{1},\ldots,M_{T}$ defined as
$M_{T}=\sum\limits_{t=1}^{T}\left[T_{d}^{t}+d(\bar{J}^{t}(\cdot),\bar{J}^{*}(\cdot))\right],$ (43)
it can be verified that the following two properties hold:
$\mathbb{E}\left[|M_{T}|\right]<\infty,$ (44)
$\mathbb{E}\left[M_{T+1}|M_{1},\ldots,M_{T}\right]=M_{T}.$ (45)
Thus, the stochastic process $M_{T}$ is a martingale. The second condition in Theorem 2 indicates that the variance of the increments of the martingale satisfies the summability condition, i.e., $\sum\limits_{t=1}^{T}t^{-2}\mathbb{V}\left\{T_{d}^{t}\right\}<\infty$. Therefore, according to the martingale convergence theorem [38], we have
[…]
Equation (46) can be equivalently rewritten as
[…]
By combining equations (34) and (47), we have
$\sup\limits_{T}\frac{1}{T}\sum\limits_{t=1}^{T}T_{d}\left(\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}\right)<0.$ (48)
According to quation (48), the following claim can be proved:
[…]
The proof of claim (49) can be performed by analyzing two possible choices of $T_{d}(\cdot)$ in equation (31), which are briefly explained as follows.
Case I:  $T_{d}(\cdot)=\log(\cdot)$. In this case, equation (48) can be rewritten as:
$\sup\limits_{T}\frac{1}{T}\sum\limits_{t=1}^{T}\log\left(\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}\right)<0.$ (50)
By choosing an appropriate $\epsilon_{L}$ as
$\epsilon_{L}=-\sup\limits_{T}\frac{1}{T}\sum\limits_{t=1}^{T}\log\left(\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}\right),$ (51)
we obtain
$\sum\limits_{t=1}^{T}\log\left(\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}\right)$ $<-T\epsilon_{L}$ (52)
$\log~{}L_{T}(\Omega)-\log~{}L_{0}(\Omega)$ $<-T\epsilon_{L}$ (53)
$L_{T}(\Omega)$ $<L_{0}(\Omega)\exp(-T\epsilon_{L}),$ (54)
which completes the proof of claim (49) for Case I.
Case II:  $T_{d}(\cdot)=\sqrt{\cdot}-1$. Equation (48) can be rewritten as:
$\sup\limits_{T}\frac{1}{T}\sum\limits_{t=1}^{T}\left(\sqrt{\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}}-1\right)<0.$ (55)
In this case, by choosing
$\epsilon_{L}=-\sup\limits_{T}\frac{2}{T}\sum\limits_{t=1}^{T}\left(\sqrt{\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}}-1\right),$ (56)
we have
$\sum\limits_{t=1}^{T}\left(\sqrt{\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}}-1\right)<\frac{1}{2}T\epsilon_{L}.$ (57)
Using the Taylor expansion on the left side of the inequality(57), the following inequality can be obtained:
$\log L_{T}(\Omega)<\log L_{0}-T\epsilon_{L},$ (58)
which completes the proof of claim (49) for Case II by taking the exponential $\exp(\cdot)$ on both sides of the inequality.
Step 2: Probability analysis. According to the iterative derivation provided in equation (23) and the probability distribution in (26), the posterior pdf can be represented as
$\mathbb{P}(\bar{J}\in\Omega)$ $=\int_{\Omega}F_{\rm pdf}^{t}(\bar{J}){\rm d}{v_{\mathcal{J}}(\bar{J})}$ (59)
$=\frac{\int_{\Omega}R_{t}F_{\rm pdf}^{0}({\rm d}v_{\mathcal{J}}(\bar{J}))}{\int_{\bar{\mathcal{J}}}R_{t}F_{\rm pdf}^{0}({\rm d}{v_{\mathcal{J}}(\bar{J})})}$ (60)
$=\frac{L_{t}(\Omega)}{\int_{\bar{\mathcal{J}}}R_{t}F_{\rm pdf}^{0}({\rm d}{v_{\mathcal{J}}(\bar{J})})}.$ (61)
Then, according to the claim (49), we have
[…]
As a result, by defining
$P_{0}=\frac{L_{0}(\Omega)}{\max\limits_{t}\int_{\bar{\mathcal{J}}}R_{t}F_{\rm pdf}^{t}({\rm d}{v_{\mathcal{J}}(\bar{J})})},$ (63)
we obtain
[…]
which completes the proof of Theorem 2.
Remark 6.
Theorem 2 characterizes the convergence of the predictive density $\bar{J}^{T}$ under the proposed learning algorithm. Specifically, as time progresses, $\bar{J}^{T}$ converges to a small neighborhood ${\mathcal{B}}(\bar{J}^{*},\delta_{\bar{J}})$. In the proof of Theorem 2, the convergence result is established by constructing a martingale process and applying the martingale convergence theorem to show that the probability of $\bar{J}^{T}$ belonging to the complement of the neighborhood of $\bar{J}^{*}$ decreases exponentially. This neighborhood is defined by $\bar{J}^{*}$ and a small positive constant $\delta_{\bar{J}}$, where $\bar{J}^{*}$ depends on the system dynamics and noises (as shown in (25)). The parameter $\delta_{\bar{J}}$ is a small positive constant that characterizes the size of the neighborhood. As indicated in (51) and (56), the parameter $\delta_{\bar{J}}$ affects the convergence rate $\epsilon_{L}$, as $\delta_{\bar{J}}$ determines the set $\Omega={\mathcal{B}}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}})$.
Remark 7.
The convergence of Theorem 2 holds for both the Hellinger distance and the KL divergence, which are two widely used methods for describing the similarity between different pdfs. On the one hand, discussing both metrics is helpful because the convergence of one metric does not necessarily guarantee the convergence of the other. On the other hand, the proof approach of the proposed method has the potential to be extended to other metrics, such as the Wasserstein distance, Jensen-Shannon divergence, and Bhattacharyya distance, etc. For these metrics, the core of the proof would lie in constructing an appropriate function $T_{d}(\cdot)$ and the corresponding martingale process $M_{T}$, and then deriving the convergence of the learned function $\bar{J}^{t}$ through the convergence of the martingale process.
IV-D Analysis of Control Regret
In Algorithm 2, the control law ${\bm{g}}_{t}$ is selected by maximizing the predictive density $\bar{J}^{t}({\bm{g}})$. However, the choice of predictive density $\bar{J}^{t}({\bm{g}})$ balances exploration and exploitation, i.e., the trade-off between exploring new control laws and exploiting the currently known control laws. The previous subsection focused on the convergence of the learning process, while this subsection focuses on the cost of the control process. Therefore, to distinguish the state under the controller ${\bm{g}}_{t}(\cdot)$ from the optimal controller ${\bm{g}}^{*}(\cdot)$, we use ${\bm{x}}^{*}(k)$ to denote the state under the controller ${\bm{g}}^{*}(\cdot)$.
To facilitate the analysis of controlling regret in the proposed method, let $\bar{J}_{t}(\cdot)\in\bar{\mathcal{J}}$ be a stochastic function with its pdf denoted as $F^{t}_{\rm pdf}(\bar{J})$. To avoid confusion, the function chosen in the $t$-th segment is $\bar{J}^{t}$, which can be regarded as a constant, while $\bar{J}_{t}(\cdot)$ is a random variable. Therefore, the function $\bar{J}^{t}$ can be considered as an instance of the random variable $\bar{J}_{t}(\cdot)$. After the $t$-th segment, the collected sampled data can be used to calculate the cost, which is denoted as $\bar{J}_{t}({\bm{g}}_{t})$ and is available in the learning process. Compared with the unknown optimal reward $\bar{J}^{*}({\bm{g}}^{*})$, the expected regret over the $t$-th period can be defined as:
$R_{e}({\bm{g}}_{t})=\left|\mathbb{E}\left[\bar{J}^{*}({\bm{g}}^{*})\right]-\mathbb{E}\left[\bar{J}_{t}({\bm{g}}_{t})\right]\right|.$ (65)
he upper bound of this regret function is provided in the following theorem.
Theorem 3.
Let Assumptions 1-3 hold. With the proposed parameterization approach in (12) and the implementation of the TS-based learning in Algorithm 2, the regret defined in (65) satisfies the following upper bound:
$R_{e}({\bm{g}}_{t})$ (66)
$\leq$ $\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))+\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}\underline{J_{m}}^{2}}$
[…]
where $\epsilon_{J}>0$ is a constant.
Proof.
According to the triangle inequality, the control regret considered in (65) can be decomposed into the sum of two terms as
$R_{e}({\bm{g}}_{t})$
$\leq$ $\left|\mathbb{E}\left[\bar{J}_{t}({\bm{g}}_{t})\right]-\mathbb{E}\left[\bar{J}^{*}({\bm{g}}_{t})\right]\right|+\left|\mathbb{E}\left[\bar{J}^{*}({\bm{g}}_{t})\right]-\mathbb{E}\left[\bar{J}^{*}({\bm{g}}^{*})\right]\right|,$ (67)
We first discuss the term $\left|\mathbb{E}\left[\bar{J}_{t}({\bm{g}}_{t})\right]-\mathbb{E}\left[\bar{J}^{*}({\bm{g}}_{t})\right]\right|$ as follows:
$\left|\mathbb{E}\left[\bar{J}_{t}({\bm{g}}_{t})\right]-\mathbb{E}\left[\bar{J}^{*}({\bm{g}}_{t})\right]\right|$ (68)
$=$ $\left|\int_{\bar{\mathcal{J}}}\bar{J}({\bm{g}}_{t})F^{t}_{\rm pdf}(\bar{J}){\rm d}v_{\bar{\mathcal{J}}}(\bar{J})-\int_{\bar{\mathcal{J}}}\bar{J}({\bm{g}}_{t})F^{*}_{\rm pdf}(\bar{J}){\rm d}v_{\bar{\mathcal{J}}}(\bar{J})\right|$
$\leq$ $\int_{\bar{\mathcal{J}}}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J})$ (69)
$\leq$ $\int_{\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J})$ (70)
$+\int_{\mathcal{B}^{c}(\bar{J}^{*},\delta_{\bar{J}})}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J}).$ (71)
Furthermore, the first term in (70) can be bounded as
$\int_{\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J})$
$\leq$ $v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\cdot\max\limits_{{\bm{g}}_{t}}\bar{J}({\bm{g}}_{t})$
$~{}\cdot\left[1-\int\limits_{\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})}F^{t}_{\rm pdf}(\bar{J}){\rm d}v_{\bar{\mathcal{J}}}(\bar{J})\right]$
$\leq$ $(1-P_{0}\exp(-t\epsilon_{L}))\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})).$ (72)
According to the boundedness of the function space $\bar{\mathcal{J}}$ and Theorem 2, the second term in (70) can be further bounded as
$\int_{\mathcal{B}^{c}(\bar{J}^{*},\delta_{\bar{J}})}\left|\left[F^{t}_{\rm pdf}(\bar{J})-F^{*}_{\rm pdf}(\bar{J})\right]\bar{J}({\bm{g}}_{t})\right|{\rm d}v_{\bar{\mathcal{J}}}(\bar{J})$
$\leq$ $2\bar{M}_{\bar{\mathcal{J}}}P_{0}\exp(-t\epsilon_{L}).$ (73)
By combining equations (Proof) and (Proof), we have
$\left|\mathbb{E}\left[\bar{J}_{t}({\bm{g}}_{t})\right]-\mathbb{E}\left[\bar{J}^{*}({\bm{g}}_{t})\right]\right|$ (74)
$\leq$ $(1-P_{0}\exp(-t\epsilon_{L}))\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))+2\bar{M}_{\bar{\mathcal{J}}}P_{0}\exp(-t\epsilon_{L})$
$\leq$ $\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\!+\!\left[2\bar{M}_{\bar{\mathcal{J}}}\!-\!\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\right]P_{0}\exp(-t\epsilon_{L}).$
Then the upper bound of the second term in equation (Proof), i.e., $\left|\mathbb{E}\left[\bar{J}^{*}({\bm{g}}_{t})\right]-\mathbb{E}\left[\bar{J}^{*}({\bm{g}}^{*})\right]\right|$, can be derived as follows:
$\left|\mathbb{E}\left[\bar{J}^{*}({\bm{g}}_{t})\right]-\mathbb{E}\left[\bar{J}^{*}({\bm{g}}^{*})\right]\right|$
$=$ $\left|\int_{\bar{\mathcal{J}}}\bar{J}^{*}({\bm{g}}_{t})F^{*}_{\rm pdf}(\bar{J}^{*}){\rm d}v_{\bar{\mathcal{J}}}(\bar{J}^{*})\!-\!\int_{\bar{\mathcal{J}}}\bar{J}^{*}({\bm{g}}^{*})F^{*}_{\rm pdf}(\bar{J}^{*}){\rm d}v_{\bar{\mathcal{J}}}(\bar{J}^{*})\right|$
$\leq$ $\int_{\bar{\mathcal{J}}}\left|\bar{J}^{*}({\bm{g}}_{t})-\bar{J}^{*}({\bm{g}}^{*})\right|F^{*}_{\rm pdf}(\bar{J}^{*}){\rm d}v_{\bar{\mathcal{J}}}(\bar{J}^{*}).$ (75)
By recalling equation (25) that $\bar{J}^{*}\propto\frac{1}{J^{*}}$ and writing the proportional coefficient as $\epsilon_{J}$, we have $\bar{J}^{*}=\frac{\epsilon_{J}}{J^{*}}$. According to Assumption 2 that the function $J({\bm{g}})$ is Lipschitz continuous and bounded, we obtain
$\left|\bar{J}^{*}({\bm{g}}_{t})-\bar{J}^{*}({\bm{g}}^{*})\right|=\left|\frac{\epsilon_{J}}{J^{*}({\bm{g}}_{t})}-\frac{\epsilon_{J}}{J^{*}({\bm{g}}^{*})}\right|$
$=$ $\epsilon_{J}\frac{\left|J^{*}({\bm{g}}_{t})-J^{*}({\bm{g}}^{*})\right|}{|J^{*}({\bm{g}}_{t})||J^{*}({\bm{g}}^{*})|}\leq\frac{\epsilon_{J}L_{J}}{\underline{J_{m}}^{2}}\|{\bm{g}}_{t}-{\bm{g}}^{*}\|_{\mathcal{L}_{2}}.$ (76)
Using the upper bound provided in Theorem 1, inequality (Proof) can be further derived as
$\left|\bar{J}^{*}({\bm{g}}_{t})-\bar{J}^{*}({\bm{g}}^{*})\right|\leq\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}\underline{J_{m}}^{2}}.$ (77)
Furthermore, the definition of probabilities lead to
$\int_{\bar{\mathcal{J}}}F^{*}_{\rm pdf}(\bar{J}^{*}){\rm d}v_{\bar{\mathcal{J}}}(\bar{J}^{*})=1.$ (78)
Then, the upper bound of the second term in equation (Proof) can be derived as
$\left|\mathbb{E}\left[\bar{J}^{*}({\bm{g}}_{t})\right]-\mathbb{E}\left[\bar{J}^{*}({\bm{g}}^{*})\right]\right|$
$\leq$ $\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}\underline{J_{m}}^{2}}.$ (79)
Combining inequalities (74) and (Proof), the regret can be bounded as
$R_{e}({\bm{g}}_{t})\leq$ $\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))$
$+\left[2\bar{M}_{\bar{\mathcal{J}}}-\bar{J_{b}}v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))\right]P_{0}\exp(-t\epsilon_{L})$
$+\frac{\epsilon_{J}L_{J}\sqrt{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}}{\sqrt{2^{n}}\underline{J_{m}}^{2}},$ (80)
which completes the proof.
Remark 8.
Theorem 3 provides an upper bound for the regret of the proposed learning Algorithm 2. The regret is decomposed into three terms. The first term represents the regret caused by the neighborhood $\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})$ used in Theorem 2. Intuitively, Theorem 2 shows that the probability of $\bar{J}^{t}\in\mathcal{B}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}})$ converges to zero but does not analyze the characteristics of the learned distribution function within the neighborhood. Therefore, the characterization of the convergence of the learned function $\bar{J}^{t}$ in Theorem 2 inevitably introduces regret related to the size of the selected neighborhood, which is a constant depending only on the parameter $\delta_{\bar{J}}$. The second term represents the regret caused by the difference between the learned controller $\bar{\bm{g}}_{t}$ and the potential optimal function ${\bm{g}}^{*}$, which is evaluated in Theorem 1 and is caused by the error introduced by the parameterization of the function space. Similar to the first term, the second term is also a constant that does not change during the learning process. The third term is the changing regret during the learning process, which is caused by the gap between the learned distribution function $F^{t}_{\rm pdf}(\cdot)$ and the true value $F^{*}_{\rm pdf}(\cdot)$. It reflects the control regret caused by the unknown system dynamics and noise.
The regret bound in Theorem 3 describes the expected gap in reward between the control law ${\bm{g}}_{t}$ adopted in the $t$-th segment of the proposed method and the potential optimal control law ${\bm{g}}^{*}$. In the literature related to TS, the average per period regret over $T$-periods is often discussed, which is defined as $R_{e}^{T}=\frac{1}{T}\left|\sum_{t=0}^{T}R_{e}({\bm{g}}_{t})\right|$. Based on the regret upper bound (66) in Theorem 3 and the analysis of each term in Remark 8, we can conclude that the average per period regret over $T$-periods for the proposed method satisfies $R_{e}^{T}=\mathcal{C}+\mathcal{O}(1/T)$, where $\mathcal{C}$ is the constant formed by the first two terms in (66), and $\mathcal{O}(1/T)$ reflects the asymptotic property of the third term in (66). Compared to existing literature [31], the proposed method not only characterizes the convergence of the learning process but also discusses the impact of function space parameterization and the hyperparameter $\delta_{\bar{J}}$ on the closed-loop regret. Specifically, for sufficiently small $\delta_{\bar{J}}$, the proposed method can converge to the optimal solution ${\bm{g}}^{*}$ within the constructed function space, which represents the unavoidable regret compared with that of selecting the control law from an infinite-dimensional function space. The regret can be further reduced by constructing the function space $\mathcal{G}$ more appropriately. It is worth noting that a function space $\mathcal{G}$ with more possibilities also leads to increased computational complexity and slower convergence.
V Conclusion
This work proposes an ALC method for unknown nonlinear systems based on the TS algorithm, aiming to optimize controller design through online learning. The proposed method constructs a Hilbert space $\mathcal{G}$ for potential candidate control laws and optimizes the selection of control laws within this space using online data. the parameterization method provides a well-structured function space for control law optimization, and the resulting potential performance degradation is proved to be bounded. Furthermore, through an active online sampling strategy, the learned cost function is proved to converge to a neighborhood of the true cost function exponentially. Additionally, the control regret during the online learning process is analyzed, including the constant regret caused by the function space construction and the exponentially convergent regret during the online learning process. Numerical experiments validate the effectiveness of the proposed method and the correctness of the related theorems. Overall, the TS method effectively balances exploration and exploitation, rapidly improving control performance in the early stages of learning. However, the exploration near the steady-state may lead to a slight decrease in control performance. Future work will focus on further improving the convergence speed and control performance in the steady-state phase of the learning process.
Acknowledgment
The authors would like to express their gratitude to Yongjie Liu and Yuchen Yang from Peking University for their valuable assistance in improving the quality of the manuscript.